### 13 Feb 2024 v1

@@ 1 Introduction
Figure 1: Instances of transductive active learning where the target space $\mathcal{A}$ is shown in blue and the sample space $\mathcal{S}$ is shown in gray. The points denote plausible observations within $\mathcal{S}$ to “learn” $\mathcal{A}$. In (A), the target space contains “everything” within $\mathcal{S}$. In (B, C, D), one makes observations “directed” towards learning about a particular target.
@@ 4 Few-Shot Fine-Tuning of Neural Networks
Using a neural network to parameterize $f$, we can approximate $f$ by a stochastic process by extracting embeddings $\boldsymbol{\phi}(\boldsymbol{x})$, e.g., latent representations, from the network. Collecting them as rows in the design matrix $\boldsymbol{\Phi}$ of inputs $X$, we approximate the network by the linear function ${\boldsymbol{f}_{\!\!X}=\boldsymbol{\Phi}\boldsymbol{\theta}}$ with weights $\boldsymbol{\theta}$. Approximating the weights by ${\boldsymbol{\theta}\sim\mathcal{N}({}\boldsymbol{\mu},\boldsymbol{\Sigma})}$ implies $\smash{\boldsymbol{f}_{\!\!X}\sim\mathcal{N}({}\boldsymbol{\Phi}\boldsymbol{\mu},\boldsymbol{\Phi}\boldsymbol{\Sigma}\boldsymbol{\Phi}^{\top})}$. The embeddings $\boldsymbol{\phi}(\boldsymbol{x})$ are commonly taken either as latent representations encoded by $f$ (Lee et al., 2018; Devlin et al., 2019) or parameter gradients (Khan et al., 2019; He et al., 2020; Holzmüller et al., 2023) in which case $\smash{\boldsymbol{\Phi}\boldsymbol{\Phi}^{\top}}$ is the (empirical) neural tangent kernel (Arora et al., 2019). [^8] We extensively discuss embeddings in Section G.2. We use this Gaussian approximation with parameter gradient embeddings and ${\boldsymbol{\Sigma}=\boldsymbol{I}}$ to evaluate the ITL decision rule, and select inputs for labeling and training of $f$. The covariance matrix $\boldsymbol{\Sigma}$ quantifies uncertainty about the parameters $\boldsymbol{\theta}$, and we study approximations in Section G.3.
@@
We show in Appendix B that the approximation error of $B^{\prime}_{n}=\boldsymbol{x}_{n,1:b}$ can be bounded in terms of the submodularity ratio of ITL (Das & Kempe, 2018). For example, if $\mathcal{S}\subseteq\mathcal{A}$, the problem from Equation 4 is submodular and the $B^{\prime}_{n}$ is a $(1-1/e)$-approximation of $B_{n}$ (Nemhauser et al., 1978; Krause & Golovin, 2014). The batch $B_{n}$, and therefore also $B^{\prime}_{n}$, is diverse and informative by design. Prior work has shown that when the problem from Equation 4 is submodular, the greedy solution $B^{\prime}_{n}$ is also competitive with a fully sequential “batchless” decision rule (Chen & Krause, 2013; Esfandiari et al., 2021). We discuss an efficient iterative implementation of Equation 5 in Section G.4.
@@ Results
Both ITL and CTL use the learned embeddings to observe points similar to $\mathcal{P}_{\!\!\mathcal{A}}$, and in particular, obtain more samples from the support of $\mathcal{P}_{\!\!\mathcal{A}}$ than Random sampling of $\mathcal{P}_{\!\mathcal{S}}$. We observe empirically that ITL obtains samples from $\mathcal{P}_{\!\!\mathcal{A}}$ at more than twice the rate of CosineSimilarity, which translates to a significant improvement in accuracy in more difficult learning tasks, while requiring fewer (labeled) samples from $\mathcal{P}_{\!\mathcal{S}}$. This phenomenon manifests for both MNIST and CIFAR-100, as well as imbalanced data distributions $\mathcal{P}_{\!\mathcal{S}}$ or imbalanced reference samples from $\mathcal{P}_{\!\!\mathcal{A}}$ (cf. Section G.6). The improvement in accuracy appears to further increase in the large-data regime, where the learning tasks become more difficult (cf. Figure 7 in Appendix G).
@@ 5 Related Work
Our work extends several previous works in active learning and experimental design (Chaloner & Verdinelli, 1995; Settles, 2009). Uncertainty sampling (Lewis & Catlett, 1994, US) is one of the most popular active learning methods. US selects points $\boldsymbol{x}$ with high prior uncertainty: ${\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\sigma_{n-1}^{2}(\boldsymbol{x})}$. This is in stark contrast to ITL which selects points $\boldsymbol{x}$ that minimize posterior uncertainty about $\mathcal{A}$. It can be seen that US is the special “undirected” (i.e., $\mathcal{S}\subseteq\mathcal{A}$) case of ITL when the observation noise is homoscedastic (cf. Section C.2).
@@
Substantial work on active learning has studied entropy-based criteria in parameter-space, most notably BALD (Houlsby et al., 2011), which selects actions ${\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{X}}\mathrm{I}\!\left(\boldsymbol{\theta};y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n-1}\right)}$, where $\boldsymbol{\theta}$ is the random parameter vector of a parametric model (e.g., via Bayesian deep learning). Such methods are inherently “undirected” in the sense that they do not facilitate learning on specific prediction targets. In contrast, ITL operates in output-space where it is straightforward to specify prediction targets, and which is computationally easier.
@@
More recently, “directed” active learning methods have been proposed for the problem of determining the optimum of an unknown function, also known as best-arm identification (Audibert et al., 2010) or pure exploration bandits (Bubeck et al., 2009). Entropy search methods (Hennig & Schuler, 2012; Hernández-Lobato et al., 2014) are widely used and select ${\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{X}}\mathrm{I}\!\left(\boldsymbol{x}^{*};y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n-1}\right)}$ in input-space where ${\boldsymbol{x}^{*}=\operatorname*{arg\,max}_{\boldsymbol{x}}f_{\boldsymbol{x}}}$. Similarly to ITL, output-space entropy search methods (Hoffman & Ghahramani, 2015; Wang & Jegelka, 2017), which select ${\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{X}}\mathrm{I}\!\left(f^{*};y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n-1}\right)}$ with ${f^{*}=\max_{\boldsymbol{x}}f_{\boldsymbol{x}}}$, are more computationally tractable. In fact, entropy search methods are a special case of transductive active learning with target space $\mathcal{A}=\{\boldsymbol{x}^{*}\}$. When subsampling target spaces via Thompson sampling (Russo et al., 2018), ITL approximates these output-space methods. Bogunovic et al. (2016) analyze TruVar in the context of Bayesian optimization and level set estimation. TruVar is akin to a variance-based variant of ITL with a similar notion of “target space”, but their algorithm and analysis rely on a threshold scheme which requires that ${\mathcal{A}\subseteq\mathcal{S}}$. Fiez et al. (2019) introduce the transductive linear bandit problem, which is a special case of transductive active learning limited to a linear function class and with the objective of determining the maximum within an initial candidate set.
@@ 6 Limitations
In this work, we focus solely on sequential decision-making given some model, rather than asking how one should construct such a model so that it is representative of the ground truth. Learning stochastic models and approximations thereof have been the subject of much recent work (Blundell et al., 2015; Maddox et al., 2019; Daxberger et al., 2021; Lin et al., 2023). Specifically, ITL benefits if the model $f$ captures the right “correlations” between points and if its “uncertainty” about the prediction at a specific point is accurate. [^11] For example, the results from Section 4 rely on embeddings being trained to capture the distribution shift.
@@ Acknowledgements
This project was supported in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and Innovation Program Grant agreement no. 815943, the Swiss National Science Foundation under NCCR Automation, grant agreement 51NF40 180545, and by a grant of the Hasler foundation (grant nr: 21039).
We thank Armin Lederer, Johannes Kirschner, Jonas Rothfuss, Lars Loch, Manish Prajapat, Nicolas Emmenegger, Parnian Kassraie, and Scott Sussex for feedback. We thank Anton Baumann for helpful discussions.
@@ A.2 Gaussian Processes
For Gaussian random vectors $\boldsymbol{f}$ and $\boldsymbol{y}$, entropy is ${\mathrm{H}\left[\boldsymbol{f}\right]=\frac{n}{2}\log(2\pi e)+\frac{1}{2}\log\left|\mathrm{Var}{}\left[\boldsymbol{f}\right]\right|}$, information gain is ${\mathrm{I}\left(\boldsymbol{f};\boldsymbol{y}\right)=\frac{1}{2}\left(\log\left|\mathrm{Var}{}\left[\boldsymbol{y}\right]\right|-\log\left|\mathrm{Var}{}\left[\boldsymbol{y}\mid\boldsymbol{f}\right]\right|\right)}$ and
@@ Appendix B Provably Efficient Non-adaptive Batch Selection and its Relation to Submodularity
$F_{n}(B)\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\mathrm{I}\!\left(\boldsymbol{f}_{\!\!\mathcal{A}};\boldsymbol{y}_{\!B}\ \middle|\ \mathcal{D}_{n-1}\right)=\mathrm{H}\!\left[\boldsymbol{f}_{\!\!\mathcal{A}}\ \middle|\ \mathcal{D}_{n-1}\right]-\mathrm{H}\!\left[\boldsymbol{f}_{\!\!\mathcal{A}}\ \middle|\ \mathcal{D}_{n-1},\boldsymbol{y}_{\!B}\right]$
is non-negative and monotone (i.e., $F_{n}(B^{\prime})\leq F_{n}(B)$ for any ${B^{\prime}\subseteq B\subseteq\mathcal{S}}$) since conditional entropy is monotone.
@@
$=\mathrm{I}\left(\boldsymbol{f}_{\!\!\mathcal{A}};y_{\boldsymbol{x}}\mid\mathcal{D}_{n-1},\boldsymbol{y}_{\!B}\right),$
@@
Corollary B.2 (Efficiency of batch selection via conditional embeddings).
For any $n,k\geq 1$, the greedy solution $B^{\prime}_{n,k}$ provides a $\left(1-e^{-\kappa_{n}(b)}\right)$-approximation of $B_{n,k}$.
The special “undirected” case
If $\mathcal{S}\subseteq\mathcal{A}$, it is well known (Srinivas et al., 2009) that $F_{n}$ is submodular, which implies that $\kappa_{n}(k)\geq 1$ for all $k\geq 1$.
@@ C.2 Undirected Case of ITL
If additionally $f$ is a GP and the observation noise is homoscedastic, then
@@
where $(i)$ follows from the monotonicity of conditional information gain since $\mathcal{S}\subseteq\mathcal{A}$ which implies $\mathcal{D}_{i}\perp y_{\boldsymbol{x}}\mid\boldsymbol{f}_{\!\!\mathcal{A}}$ for all $\boldsymbol{x}\in\mathcal{S}$; $(ii)$ uses the objective of ITL; $(iii)$ uses the chain rule of information gain; $(iv)$ uses ${\boldsymbol{y}_{\!X}\perp\boldsymbol{f}_{\!\!\mathcal{A}}\mid\boldsymbol{f}_{\!\!X}}$ and the data processing inequality. The conditional independences follow from the assumption that the observation noise is independent. ∎

### 12 Mar 2024 v2

@@ 1 Introduction
Figure 1: Instances of transductive active learning where the target space $\mathcal{A}$ is shown in blue and the sample space $\mathcal{S}$ is shown in gray. The points denote plausible observations within $\mathcal{S}$ to “learn” $\mathcal{A}$. In (A), the target space contains “everything” within $\mathcal{S}$ as well as points outside $\mathcal{S}$. In (B, C, D), one makes observations directed towards learning about a particular target. Prior work on active learning has focused on the instance $\mathcal{A}=\mathcal{S}$.
@@ 4 Few-Shot Fine-Tuning of Neural Networks
The approximation of NNs $f(\cdot;\boldsymbol{\theta})$ as linear functions in a latent embedding space,
$f(\boldsymbol{x};\boldsymbol{\theta})\approx\boldsymbol{\beta}^{\top}\boldsymbol{\phi}_{\boldsymbol{\theta}}(\boldsymbol{x}),$ (3)
with weights $\boldsymbol{\beta}\in\mathbb{R}^{p}$ and embeddings $\boldsymbol{\phi}_{\boldsymbol{\theta}}:\mathcal{X}\to\mathbb{R}^{p}$ has been studied by a rich line of prior work. Common choices of embeddings include last-layer embeddings (Devlin et al., 2019; Holzmüller et al., 2023), neural tangent embeddings arising from neural tangent kernels (Jacot et al., 2018) which are motivated by their relationship to the training and fine-tuning of ultra-wide NNs (Arora et al., 2019; Lee et al., 2019; Khan et al., 2019; He et al., 2020; Malladi et al., 2023), and loss gradient embeddings (Ash et al., 2020). We provide a comprehensive overview of embeddings in Section G.2. Now, supposing the prior $\boldsymbol{\beta}\sim\mathcal{N}({}\boldsymbol{0},\boldsymbol{\Sigma})$, often with ${\boldsymbol{\Sigma}=\boldsymbol{I}}$ (Khan et al., 2019; He et al., 2020; Antorán et al., 2022; Wei et al., 2022), the approximation of $f$ from Equation 3 is a Gaussian process with kernel $k(\boldsymbol{x},\boldsymbol{x^{\prime}})=\boldsymbol{\phi}_{\boldsymbol{\theta}}(\boldsymbol{x})^{\top}\boldsymbol{\Sigma}\boldsymbol{\phi}_{\boldsymbol{\theta}}(\boldsymbol{x^{\prime}})$. The covariance matrix $\boldsymbol{\Sigma}$ quantifies uncertainty about the weights $\boldsymbol{\beta}$, and we study approximations in Section G.3. In our experiments, we use this Gaussian approximation with loss gradient embeddings [^8] We observe essentially the same performance with neural tangent embeddings, cf. Section G.2. and ${\boldsymbol{\Sigma}=\boldsymbol{I}}$ to evaluate the ITL decision rule, and select inputs for labeling and training $f$. Notably, the approximation of Equation 3 is only used to guide the active data selection, and not for inference.
Whereas prior work has studied the generalization error when using the posterior mean of $\boldsymbol{\beta}$ for prediction at a test point (Wei et al., 2022), our results from Section 3 bound the epistemic posterior uncertainty about this prediction given that the model is trained using data selected by ITL.
@@
We show in Appendix B that the approximation error of $B^{\prime}_{n}=\boldsymbol{x}_{n,1:b}$ can be bounded in terms of the submodularity ratio of ITL (Das & Kempe, 2018). For example, if $\mathcal{S}\subseteq\mathcal{A}$, the problem from Equation 5 is submodular and the $B^{\prime}_{n}$ is a $(1-1/e)$-approximation of $B_{n}$ (Nemhauser et al., 1978; Krause & Golovin, 2014). The batch $B_{n}$, and therefore also $B^{\prime}_{n}$, is diverse and informative by design. We discuss an efficient iterative implementation of Equation 6 in Section G.4. Prior work has shown that the greedy solution $B^{\prime}_{n}$ is also competitive with a fully sequential “batchless” decision rule (Chen & Krause, 2013; Esfandiari et al., 2021). [^9] They prove this for the case where Equation 5 is submodular, but their results readily generalize to “approximate” submodularity.
@@ Results
Both ITL and CTL use the learned embeddings to observe points similar to $\mathcal{P}_{\!\!\mathcal{A}}$, and in particular, obtain more samples from the support of $\mathcal{P}_{\!\!\mathcal{A}}$ than Random sampling of $\mathcal{P}_{\!\mathcal{S}}$. We observe empirically that ITL obtains samples from $\mathcal{P}_{\!\!\mathcal{A}}$ at more than twice the rate of CosineSimilarity, which translates to a significant improvement in accuracy in more difficult learning tasks, while requiring fewer (labeled) samples from $\mathcal{P}_{\!\mathcal{S}}$. This phenomenon manifests for both MNIST and CIFAR-100, as well as imbalanced data distributions $\mathcal{P}_{\!\mathcal{S}}$ or imbalanced reference samples from $\mathcal{P}_{\!\!\mathcal{A}}$ (cf. Section G.6). The improvement in accuracy appears to further increase in the large-data regime, where the learning tasks become more difficult (cf. Figure 7 in Appendix G). Akin to the previously identified scaling trend with size of the pre-training dataset (Tamkin et al., 2022), this suggests a potential scaling trend where the improvement of ITL over random batch selection grows as models are fine-tuned on a larger pool of data.
@@ 5 Related Work
Our work extends several previous works in active learning and experimental design (Chaloner & Verdinelli, 1995; Settles, 2009). Uncertainty sampling (US, Lewis & Catlett, 1994) is one of the most popular active learning methods. US selects points $\boldsymbol{x}$ with high prior uncertainty: ${\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\sigma_{n-1}^{2}(\boldsymbol{x})}$. This is in stark contrast to ITL which selects points $\boldsymbol{x}$ that minimize posterior uncertainty about $\mathcal{A}$. It can be seen that US is the special “undirected” (i.e., $\mathcal{S}\subseteq\mathcal{A}$) case of ITL when the observation noise is homoscedastic (cf. Section C.2).
@@
Substantial work on active learning has studied entropy-based criteria in parameter-space, most notably BALD (Houlsby et al., 2011; Kirsch et al., 2019), which selects actions ${\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{X}}\mathrm{I}\!\left(\boldsymbol{\theta};y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n-1}\right)}$, where $\boldsymbol{\theta}$ is the random parameter vector of a parametric model (e.g., via Bayesian deep learning). Such methods are inherently “undirected” in the sense that they do not facilitate learning on specific prediction targets. In contrast, ITL operates in output-space where it is straightforward to specify prediction targets, and which is computationally easier.
@@
More recently, “directed” active learning methods have been proposed for the problem of determining the optimum of an unknown function, also known as best-arm identification (Audibert et al., 2010) or pure exploration bandits (Bubeck et al., 2009). Entropy search methods (Hennig & Schuler, 2012; Hernández-Lobato et al., 2014) are widely used and select ${\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{X}}\mathrm{I}\!\left(\boldsymbol{x}^{*};y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n-1}\right)}$ in input-space where ${\boldsymbol{x}^{*}=\operatorname*{arg\,max}_{\boldsymbol{x}}f_{\boldsymbol{x}}}$. Similarly to ITL, output-space entropy search methods (Hoffman & Ghahramani, 2015; Wang & Jegelka, 2017), which select ${\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{X}}\mathrm{I}\!\left(f^{*};y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n-1}\right)}$ with ${f^{*}=\max_{\boldsymbol{x}}f_{\boldsymbol{x}}}$, are more computationally tractable. In fact, output-space entropy search is a special case of ITL with a stochastic target space (Hübotter et al., 2024). When subsampling target spaces via Thompson sampling (Russo et al., 2018), ITL approximates these output-space methods. Bogunovic et al. (2016) analyze TruVar in the context of Bayesian optimization and level set estimation. TruVar is akin to a variance-based variant of ITL with a similar notion of “target space”, but their algorithm and analysis rely on a threshold scheme which requires that ${\mathcal{A}\subseteq\mathcal{S}}$. Fiez et al. (2019) introduce the transductive linear bandit problem, which is a special case of transductive active learning limited to a linear function class and with the objective of determining the maximum within an initial candidate set.
@@ 6 Limitations
In this work, we focus solely on sequential decision-making given some model, rather than asking how one should construct such a model so that it is representative of the ground truth. Learning stochastic models and approximations thereof have been the subject of much recent work (Blundell et al., 2015; Maddox et al., 2019; Daxberger et al., 2021; Antorán et al., 2022; Lin et al., 2023). Specifically, ITL benefits if the model $f$ captures the right “correlations” between points and if its “uncertainty” about the prediction at a specific point is accurate. [^12] For example, the results from Section 4 rely on embeddings being trained to capture the distribution shift.
@@ Acknowledgements
Many thanks to Armin Lederer, Johannes Kirschner, Jonas Rothfuss, Lars Loch, Manish Prajapat, Nicolas Emmenegger, Parnian Kassraie, and Scott Sussex for their insightful feedback on different versions of this manuscript, as well as Anton Baumann for helpful discussions.
This project was supported in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and Innovation Program Grant agreement no. 815943, the Swiss National Science Foundation under NCCR Automation, grant agreement 51NF40 180545, and by a grant of the Hasler foundation (grant no. 21039). Jonas Hübotter was supported in part by the German Academic Scholarship Foundation (Studienstiftung).
@@ A.2 Gaussian Processes
For Gaussian random vectors $\boldsymbol{f}$ and $\boldsymbol{y}$, the entropy is ${\mathrm{H}\left[\boldsymbol{f}\right]=\frac{n}{2}\log(2\pi e)+\frac{1}{2}\log\left|\mathrm{Var}{}\left[\boldsymbol{f}\right]\right|}$, the information gain is ${\mathrm{I}\left(\boldsymbol{f};\boldsymbol{y}\right)=\frac{1}{2}\left(\log\left|\mathrm{Var}{}\left[\boldsymbol{y}\right]\right|-\log\left|\mathrm{Var}{}\left[\boldsymbol{y}\mid\boldsymbol{f}\right]\right|\right)}$, and
@@ Appendix B Provably Efficient Non-adaptive Batch Selection and its Relation to Submodularity
$F_{n}(B)$ =$\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\mathrm{I}\!\left(\boldsymbol{f}_{\!\!\mathcal{A}};\boldsymbol{y}_{\!B}\ \middle|\ \mathcal{D}_{n-1}\right)=\mathrm{H}\!\left[\boldsymbol{f}_{\!\!\mathcal{A}}\ \middle|\ \mathcal{D}_{n-1}\right]-\mathrm{H}\!\left[\boldsymbol{f}_{\!\!\mathcal{A}}\ \middle|\ \mathcal{D}_{n-1},\boldsymbol{y}_{\!B}\right]$
is non-negative and monotone, [^15] Formally, ${F_{n}(B)\geq 0}$ and ${F_{n}(B^{\prime})\leq F_{n}(B)}$ for any ${B^{\prime}\subseteq B\subseteq\mathcal{S}}$. since conditional entropy is monotone (which is also called the “information never hurts” principle).
@@
$=\mathrm{I}\left(\boldsymbol{f}_{\!\!\mathcal{A}};y_{\boldsymbol{x}}\mid\mathcal{D}_{n-1},\boldsymbol{y}_{\!B}\right)$
@@
Theorem B.2 (Efficiency of batch selection via conditional embeddings).
For any $n,k\geq 1$, the greedy solution $B^{\prime}_{n,k}$ provides a $\left(1-e^{-\kappa_{n}(k)}\right)$-approximation of $B_{n,k}$.
If $\mathcal{S}\subseteq\mathcal{A}$, it is well known (e.g., Srinivas et al., 2009) that $F_{n}$ is submodular, which implies that $\kappa_{n}(k)\geq 1$ for all $k\geq 1$.
@@ C.2 Undirected Case of ITL
If additionally $f$ is a GP then
@@
${=}\frac{1}{n}\sum_{i=0}^{n-1}\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{A}};y_{\boldsymbol{x}_{i+1}}\mid\boldsymbol{y}_{\boldsymbol{x}_{1:i}})$
@@
where $(i)$ follows from the monotonicity of conditional information gain since $\mathcal{S}\subseteq\mathcal{A}$ which implies $\mathcal{D}_{i}\perp y_{\boldsymbol{x}}\mid\boldsymbol{f}_{\!\!\mathcal{A}}$ for all $\boldsymbol{x}\in\mathcal{S}$; $(ii)$ uses the objective of ITL; $(iii)$ uses that the posterior variance of Gaussians is independent of the realization and only depends on the location of observations; $(iv)$ uses the chain rule of information gain; $(v)$ uses ${\boldsymbol{y}_{\!X}\perp\boldsymbol{f}_{\!\!\mathcal{A}}\mid\boldsymbol{f}_{\!\!X}}$ and the data processing inequality. The conditional independences follow from the assumption that the observation noise is independent. ∎